# Semi-supervised learning for White Blood Cell Semantic Segmentation

### 22 Jan 2024 v2

@@ Abstract
AI’s application in healthcare, especially in white blood cell cancer diagnosis, is hindered by two primary challenges: the lack of large-scale labeled datasets for white blood cell (WBC) segmentation and outdated segmentation methods. To address the first challenge, we must have a semi-supervised learning framework so that we can efficiently annotate that large dataset. In this work, we address this issue and propose a novel self-training pipeline with the incorporation of FixMatch. We conduct experiments on white blood cell datasets and evaluate the strengths and weaknesses of our pipeline.
@@ 1 Introduction
Cancer stands among the deadliest diseases in the world. In Vietnam and the U.S., cancer is only second to heart disease in terms of mortality rate. In 2020, it is estimated that nearly 19 million cases of cancer arise worldwide, with the number of deaths standing at approximately 10 million [1]. Cancer of the white blood cells (WBC), namely leukemia and lymphoma, are two of the most common types of blood cancers [2]. However, certain screening and diagnosis of blood cancers, particularly cancer of the white blood cells (WBC), still employ a manual process, such as in peripheral blood smear analysis. These tests often serve the purpose of classifying and counting the number of WBCs in the blood, and if the number is abnormally high or low, further examination is needed to decide if the patient is inflicted with cancer or not.
White blood cell segmentation is a rarely touched area among AI experts. In our experience, it is hard to find a dataset that both provides images and their ground truth masks to perform semantic segmentation. Most datasets concern themselves with white blood cell classification and do not contain the masks needed for segmentation. And when we found datasets that suited our purposes, the number of data was very insignificant.
Recognizing that there is a shortage of meaningful datasets in the field of white blood cell segmentation, we aim to provide a semi-supervised learning technique so that in the presence of a large-scale dataset, we can annotate that dataset most efficiently.
Our contributions to this paper are as follows:
    We aim to provide a semi-supervised learning approach to annotate a large dataset of white blood cells.
    From that semi-supervised learning approach, we should devise a more accurate, fine-tuned way to segment white blood cells.
@@ 2 Related works
2.0.1 Semi-supervised learning.
is a paradigm in which the dataset used for training is not fully annotated. It sits between supervised learning, where the whole dataset is labeled, and unsupervised learning, where no data is labeled. In the early days, semi-supervised learning was confined to traditional machine-learning approaches [5], such as semi-supervised support vector machines, graph-based, and co-training. Semi-supervised learning can be categorized into 5 groups: generative, consistency regularization, graph-based, pseudo-labeling, and hybrid methods.
@@ 2.0.2 Semi-supervised semantic segmentation
addresses the challenges of applying semi-supervised learning principles to the task of semantic image segmentation. It is a more realistic application of semi-supervised learning since semantic segmentation involves pixel-level labeling which is a more intensive labeling requirement than classification or detection. The techniques in this field can also be categorized into 5 categories: Adversarial, contrastive learning, consistency regularization, pseudo-labeling, and hybrid methods.
2.0.3 Consistency regularization
The consistency regularization method commonly used in semi-supervised semantic segmentation can be categorized into 4 main groups: Input perturbations, feature perturbations, network perturbations, and the three combined methods. The common goal for all of these sub-groups is to make the model perform consistently across all transformations applied to the unlabelled data.
2.0.4 Mean teacher.
Almost all of the methods in consistency regularization are based on the Mean Teacher model. The Mean Teacher model consists of two main components: the student model and the teacher model. The student model is the main model that is being trained on the available labeled data. Like in standard supervised learning, the student model learns to make predictions based on the labeled examples it sees. For the Teacher model, this is a secondary model whose weights are an exponential moving average (EMA) of the student model’s weights over time. The teacher model is not trained directly with gradient descent. Instead, its weights are updated as the moving average of the student model’s weights after each training step.
@@ 3 Methodology
3.1 Problem definition
Semi-supervised semantic segmentation can be formulated as follows: Let $X=\{X_{L},X_{U}\}$ be the dataset, and $X_{L}=\{x_{i}\}_{i=1}^{L}$ be the labeled subset, $Y_{L}=(y_{1},y_{2},\ldots,y_{L})$ is the labels for such set and $X_{U}=\{(x_{i})\}_{i=1}^{U}$ be the unlabeled subset. The assumption is that $L\ll U$. The objective is to minimize the following loss function:
$\min_{\theta}\sum_{x\in X_{L},y\in Y_{L}}\mathcal{L}_{s}(x,y,\theta)+\alpha\sum_{x\in X_{U}}\mathcal{L}_{u}(x,\theta)+\beta\sum_{x\in X}\mathcal{R}(x,\theta)$
$\mathcal{L}_{s}$ denotes the supervised losses with labeled data, $\mathcal{L}_{u}$ is the unsupervised loss and $\mathcal{R}$ is the regularization term. Depending on the technique chosen, this regularization term can be consistency regularization or another designed regularization technique. $\theta$ is the model’s parameters and $\alpha$ and $\beta$ are trade-off terms respectively.
@@ 3.2 Self-training (ST)
Algorithm 1 ST Pseudocode
1:Labeled training set $D_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$, Unlabeled training set $D_{U}=\{u_{i}\}_{i=1}^{U}$
2:Fully trained model $S$
3:Initialize teacher model $T$ with random weights
4:Pre-train $T$ on $D_{L}$ using supervised learning
5:Use model $T$ to infer labels on $D_{U}$ to get $\hat{D}_{U}$
6:Combine $D_{L}$ and $\hat{D}_{U}$ to create an augmented training set $D_{aug}$
7:Re-train student model $S$ on $D_{aug}$
8:return $S$
3.3 Self-training++ (ST++)
Despite the simplicity and impressive results of Self-training, it doesn’t have a mechanism to take into account the pseudo-masks that are extremely incorrect. By re-training on pseudo-masks that are extremely wrong, called unreliable samples, the model’s performance inevitably deteriorates. In contrast, the pseudo-masks that are close to the actual ground truth of the unlabeled image, called reliable samples, play a key role in improving the model’s performance. Therefore, in [16], the authors propose a way to harness the power of reliable images to compensate for unreliable ones. Instead of re-training directly on labeled data and all unlabelled data in the final step, the authors first train the model on labeled data and reliable samples. Then, they use this model to generate pseudo-masks for the unreliable data. Finally, they re-train the model on labeled data, reliable data, and unreliable data.
Same as the strong baseline the authors in [16] for plain self-training, the authors apply strong data augmentation on the unlabeled images whenever re-training on both labeled and unlabeled data.
Algorithm 2 ST++ Pseudocode
1:Labeled training set $D_{L}$, Unlabeled training set $D_{U}$
2:Fully trained student model $S$
3:Train teacher model $T$ on $D_{L}$
4:Obtain pseudo labels for $D_{U}$ using $T$
5:Select a subset of $D_{U}$ with the most reliable pseudo labels
6:Train student model $S$ on $D_{L}$ and the reliable subset of $D_{U}$
7:Re-label the remaining subset of $D_{U}$ using $S$
8:Re-train $S$ on $D_{L}$ and the entire $D_{U}$
9:return $S$
3.4 FixMatch
Inspired by the input perturbation method in the area of consistency-regularization applied in semi-supervised semantic segmentation, FixMatch [17] is a method introduced in 2020 based on the assumption that a model should produce the same prediction for an unlabeled image, regardless of whether it has been weakly or strongly augmented. At the time of its introduction, the FixMatch method outperformed all other methods in datasets such as CIFAR-100, SVHN, and STL-100 [17] in the field of image classification.
3.5 ST++ with FixMatch
Realizing that the ST++ approach has not yet fully utilized consistency regularization approaches, we aspire to incorporate FixMatch among the steps described in the pseudo-code for ST++. Realizing that the first step of training in a fully supervised manner in ST++ can be easily adjusted to incorporate FixMatch. We have done just that.
Algorithm 3 ST++ with FixMatch
1:Labeled dataset $D_{L}=\{(x_{i},y_{i})\}$, Unlabeled dataset $D_{U}=\{u_{j}\}$, Model parameters $\theta$, Confidence threshold $\tau$, Strong augmentation function $S(\cdot)$, Weak augmentation function $W(\cdot)$
2:Initialize: Train model on $L$ for a few epochs
@@ 3.5 ST++ with FixMatch
6:         Apply weak augmentation: $x^{\prime}_{i}=\alpha(x_{i})$
@@
14:         Train model on $\{(u^{\prime\prime}_{j},\hat{y}_{j})\}$ for samples with pseudo-labels, using consistency loss
@@ 4.1 Dataset
4.1.1 Dataset 1.
[14] Sourced from Jiangxi Tecom Science Corporation in China, comprises 300 images of white blood cells (WBCs), each measuring 120×120 pixels with a color depth of 24 bits. These images were captured using a Motic Moticam Pro 252A camera attached to an N800-D motorized auto-focus microscope. The blood smear samples in this dataset were stained using a novel hematology reagent specifically designed for rapid WBC staining, resulting in images that predominantly exhibit a yellow background.
Figure 1: Dataset1: A sample
4.1.2 Dataset 2
[14] Dataset 2 is composed of 100 color images, each with dimensions of 300×300 pixels, acquired from the CellaVision blog. These images predominantly feature cells in a purple hue and often include numerous red blood cells surrounding the white blood cells. This characteristic sets Dataset 2 apart from Dataset 1.
4.1.3 LISC.
[15] The LISC dataset consists of hematological images taken from the peripheral blood of healthy subjects. The dataset contains images of normal leukocytes, including basophils, eosinophils, lymphocytes, monocytes, and neutrophils, manually segmented and classified by experts. The images, collected from 8 normal subjects and amounting to 400 samples from 100 microscope slides, are in BMP format with a resolution of 720×576 pixels.
@@ 4.2 Settings
The labeled set accounts for 1/4 of the size of the training set. This proportion is in accordance with the setting used in the semi-supervised learning setting, such as ST++ [16].
The experiment was conducted on the RTX3090. The experiment was run under two settings: With FixMatch and without FixMatch. Under each of these settings, we employ three frameworks in semi-supervised learning:
  • SupOnly: The model is trained on the labeled dataset only.
  • ST: The model is trained in a standard self-training manner.
  • ST++: The model is trained in a self-traning++ framework according to the pseudo-code 2.
Without consistency, the batch size for Dataset 1, Dataset 2, and LISC was 16. With the consistency incorporated, the batch size for the labeled and the unlabeled data is 16. For Dataset 2 and LISC, the labeled data’s batch size is 16 and the unlabeled data’s batch size is 8. The network architectures are DeepLabV3+, DeepLabv2, and PSPNet. Each network architecture consists of the ResNet-50 and ResNet-101 backbone, but DeepLabv2 only has ResNet-101 backbone.
We will also run two algorithms: Self-Training (ST) and Self-Training++ (ST++)
@@ 5 Results
Model Backbone Backbone Dataset
Dataset 1 Dataset2 LISC
@@
without consistency 89.49 90.71 90.02 86.19 86.43 87.34 75.2 76.67 77.64
DeepLabv3 ResNet-50 with consistency 87.81 90.04 90.69 84.19 87.02 87.37 73.96 76.48 76.49
without consistency 88.86 89.83 89.99 86.03 86.35 87.22 75.95 75.87 76.55
DeepLabv3 ResNet-101 with consistency 86.61 89.8 90.29 85.15 87.28 87.09 75.09 77.02 76.6
without consistency 82.54 84.21 85.01 74.46 79.7 78.74 72.2 75.27 74.84
DeepLabv2 ResNet-101 with consistency 80.08 85.05 85.07 71.47 79.32 80.38 70.47 74.71 75.75
without consistency 80.27 83.08 83.64 75.76 80.2 80.92 71.17 74.69 75.12
PSPNet ResNet-50 with consistency 79.02 83.44 83.76 74.95 80.31 81.79 71.53 75.5 75.39
without consistency 81.56 83.93 84.3 77.76 80.39 81.35 72.91 75.19 75.17
PSPNet ResNet-101 with consistency 80.69 84.56 84.58 77.33 79.8 80.11 71.35 74.96 74.82
From the above results 1, there can be two conclusions:
  • There is a performance improvement when transitioning from Supervised-Only (SupOnly) to Self-Training (ST) and further to ST++, with ST++ demonstrating the highest level of performance.
  • With FixMatch incorporated, the performance on ST and ST++ framework improves in certain scenarios.
6 Discussion
When incorporating FixMatch into the supervised training stage, there was a noticeable decrease in performance on the labeled images. This is a commonly seen trade-off when we try to do two tasks in 1 stage.
Table 2: Noticeable decrease when incorporating FixMatch into the model.
Dataset1 Dataset2 LISC
@@ 6.1 Performance on the supervised training stage when incorporating FixMatch
DeepLabv3+ - ResNet-50 89.49 87.81 86.19 84.19 75.2 73.96
DeepLabv3+ - ResNet-101 88.86 86.61 86.03 85.15 75.95 75.09
@@
However, the performance in the later stage remains the same or even improves, suggesting that the trade-off in the first stage is sometimes necessary to improve the performance when extending the labeled images with unlabeled ones.
@@ 6.2 The quality of the pseudo-masks produced when incorporating FixMatch
The quality of the pseudo-masks between the two settings of FixMatch and non-FixMatch approaches varies. On the one hand, the pseudo-masks of easy, reliable images remain essentially the same between the two settings.
Figure 2: Sample of Ground truth mask, Non-FixMatch, and FixMatch mask generated from the LISC dataset.
However, for more complicated, unreliable images, the masks generated can have a large deviation from the ground truth mask.
@@
This suggests that incorporating FixMatch into the model has yet to be able to take on the nuances presented in challenging images.

### 23 Feb 2024 v3

# Semi-Supervised Semantic Segmentation using Redesigned Self-Training for White Blood Cells

@@ Abstract
Artificial Intelligence (AI) in healthcare, especially in white blood cell cancer diagnosis, is hindered by two primary challenges: the lack of large-scale labeled datasets for white blood cell (WBC) segmentation and outdated segmentation methods. These challenges inhibit the development of more accurate and modern techniques to diagnose cancer relating to white blood cells. To address the first challenge, a semi-supervised learning framework should be devised to efficiently capitalize on the scarcity of the dataset available. In this work, we address this issue by proposing a novel self-training pipeline with the incorporation of FixMatch. Self-training is a technique that utilizes the model trained on labeled data to generate pseudo-labels for the unlabeled data and then re-train on both of them. FixMatch is a consistency-regularization algorithm to enforce the model’s robustness against variations in the input image. We discover that by incorporating FixMatch in the self-training pipeline, the performance improves in the majority of cases. Our performance achieved the best performance with the self-training scheme with consistency on DeepLab-V3 architecture and ResNet-50, reaching 90.69%, 87.37%, and 76.49% on Zheng 1, Zheng 2, and LISC datasets, respectively.
@@ 1 Introduction
Cancer of the white blood cells (WBC), namely leukemia and lymphoma, are two of the most common types of blood cancers [1]. Certain screening and diagnosis of blood cancers, particularly cancer of the white blood cells (WBC), are still done manually, as in peripheral blood smear analysis. These tests frequently serve the purpose of classifying and counting the number of WBCs. Due to the time-consuming and experts required for labeling, most blood cell datasets lack high-quality annotation. Small-scale public datasets are not sufficient to train robust deep learning to be employed in real-world scenarios.
We intend to make use of a semi-supervised learning technique as we are aware that the field of white blood cell segmentation lacks meaningful datasets. Even with a limited amount of labeled data, we can still construct a robust model by effectively utilizing unlabeled datasets. In this work, we aim to provide a self-training approach by leveraging unlabeled data to boost the performance of the white blood cell segmentation. Secondly, inspired by FixMatch[7], we utilize a mechanism of consistency regularization to aid the self-training framework. Thirdly, we ran extensive experiments to measure the performance of the self-training approach for semi-supervised learning under various settings.
@@ 2 Related works
2.0.1 Medical Image Segmentation.
In recent years, with the growth of deep learning, medical image analysis has attracted a lot of advancement by automatically diagnosing cancer through radiology[2, 3], microscopy[4], etc. Noticeably, U-Net[5] is an autoencoder-based architecture with skip connections to incorporate feature maps from the encoder and decoder. It has been used extensively in segmentation. The combination of U-net with the backbone as Resnet-family[6] has shown impressive performance across many medical segmentation tasks and modalities and has become one of the most common baselines for semantic segmentation problems.
@@ 2.0.2 Semi-supervised semantic segmentation
This technique sits between supervised learning, where the whole dataset is labeled, and unsupervised learning, where no data is labeled. Through many years of development, Semi-supervised learning semantic segmentation can be categorized into two main directions: consistency regularization[7, 8], and pseudo-labeling[11]. The self-training-based approach is one of the most promising approaches due to its effective strategy to leverage unlabeled sets. Self-training ST++ [11] is a sophisticated framework that performs selective re-training by prioritizing reliable images based on holistic prediction-level stability throughout the training course.
@@ 3.1 Self-training (ST)
Figure 1: Our proposed semi-supervised semantic segmentation framework. Consistency regularization is incorporated in the first stage of training to boost the capabilities of performing towards unlabeled images with weak-to-strong mechanisms. This incorporation can be done in both the ST and ST++ framework. In this diagram, ST++ is taken as an example
3.2 Self-training with Consistency Regularization
In [11], the authors propose a better self-training scheme, called Self-Training++ (ST++). The innovation in this scheme is that instead of training on all unlabeled images at the same time, the authors divide this stage into two stages by separating unlabeled images into reliable and unreliable ones. Realizing that the ST and ST++ approaches have not yet fully utilized consistency regularization approaches commonly used in semi-supervised learning, we aspire to incorporate FixMatch[7] in the first supervised training step to train end-to-end framework with both supervised and unsupervised losses. In doing so, we allow the first supervised training stage to take advantage of both labeled and unlabeled data, rather than just labeled data in the original framework. What’s more, the weak-to-strong consistency regularization mechanism allows the model to be more robust against strong variations in the unlabeled data.
Algorithm 1 Supervised training stage in ST and ST++ with FixMatch
1:Labeled dataset $D_{L}=\{(x_{i},y_{i})\}$, Unlabeled dataset $D_{U}=\{u_{j}\}$, Model $M$ parameters $\theta$, Confidence threshold $\tau$, Strong augmentation function $S(\cdot)$, Weak augmentation function $W(\cdot)$
2:Initialize: Train model $M$ on $L$ for a few epochs
@@ 3.2 Self-training with Consistency Regularization
6:         Apply weak augmentation: $x^{\prime}_{i}=W(x_{i})$
@@
14:         Train model on $\{(u^{\prime\prime}_{j},\hat{y}_{j})\}$ using consistency loss
@@ 4.1 Dataset
Zheng 1 [9] comprises 300 images of white blood cells (WBCs), each measuring 120×120 pixels with a color depth of 24 bits. The images in this dataset predominantly exhibit a yellow background. Zheng 2 [9] is composed of 100 color images, each with dimensions of 300×300 pixels, acquired from the CellaVision blog. The LISC dataset [10] consists of hematological images taken from the peripheral blood of healthy subjects. It was collected from 8 subjects and amounting to 400 samples from 100 slides with a size of 720×576 pixels.
@@ 4.2 Settings
The experiment was run under two settings: With FixMatch and without FixMatch. Under each of these settings, we employ three frameworks in semi-supervised learning: SupOnly (trained on the labeled dataset only), ST (trained in a standard self-training manner), and ST++(trained in a self-traning++ framework). The labeled set accounts for 1/4 of the size of the training set. This proportion is in accordance with the setting used in the semi-supervised learning setting, such as ST++ [11].
Without consistency, the batch size for datasets Zheng 1, Zheng 2, and LISC is 16. With the consistency incorporated, the batch size for the labeled and the unlabeled data of Zheng 1 is 16. For Zheng 2 and LISC, the labeled data’s batch size is 16 and the unlabeled data’s batch size is 8. The network architectures are DeepLabV3+, DeepLabv2, and PSPNet. The choice of the DeepLab networks was because of their atrous convolution which enables wider context absorption. PSPNet also aggregates context at different scales. All of these attributes enable the networks to segment the nucleus and cytoplasm of the cells effectively. Each network architecture consists of the ResNet-50 and ResNet-101 backbone, but DeepLabv2 only has the ResNet-101 backbone. The learning rate for three datasets is 0.009 and follows a poly decay schedule $lr=baselr\times\left(1-\frac{iter}{total\_iter}\right)^{0.9}.$
@@ 5 Results
Model Backbone Mode Dataset
Zheng 1 Zheng 2 LISC
@@
without FixMatch 89.49 90.71 90.02 86.19 86.43 87.34 75.2 76.67 77.64
DeepLabv3 ResNet-50 with FixMatch 87.81 90.04 90.69 84.19 87.02 87.37 73.96 76.48 76.49
without FixMatch 88.86 89.83 89.99 86.03 86.35 87.22 75.95 75.87 76.55
DeepLabv3 ResNet-101 with FixMatch 86.61 89.8 90.29 85.15 87.28 87.09 75.09 77.02 76.6
without FixMatch 82.54 84.21 85.01 74.46 79.7 78.74 72.2 75.27 74.84
DeepLabv2 ResNet-101 with FixMatch 80.08 85.05 85.07 71.47 79.32 80.38 70.47 74.71 75.75
without FixMatch 80.27 83.08 83.64 75.76 80.2 80.92 71.17 74.69 75.12
PSPNet ResNet-50 with FixMatch 79.02 83.44 83.76 74.95 80.31 81.79 71.53 75.5 75.39
without FixMatch 81.56 83.93 84.3 77.76 80.39 81.35 72.91 75.19 75.17
PSPNet ResNet-101 with FixMatch 80.69 84.56 84.58 77.33 79.8 80.11 71.35 74.96 74.82
5.1 Performance across SupOnly, ST, ST++ with and without FixMatch incorporated
As described in Table 1, there is a performance improvement when transitioning from Supervised-Only (SupOnly) to Self-Training (ST) and further to ST++, with ST++ demonstrating the highest level of performance. This is not surprising, since each framework is an improvement upon one another. With SupOnly, the model only takes advantage of labeled data, while with ST and ST++, the model also harnesses the power of unlabeled data. Furthermore, ST++ has more discriminative power over ST, since it separates the unlabeled data into reliable and unreliable ones.
Table 2: Noticeable decrease when incorporating FixMatch into the model. The left and right column denotes the absence and presence of FixMatch respectively
Model + Backbone Zheng 1 Zheng 2 LISC
@@ 5.1 Performance across SupOnly, ST, ST++ with and without FixMatch incorporated
DeepLabv3+ - ResNet50 89.49 87.81 86.19 84.19 75.2 73.96
DeepLabv3+ - ResNet101 88.86 86.61 86.03 85.15 75.95 75.09
@@
With FixMatch incorporated, the performance of the ST and ST++ framework improves in the majority of cases, indicating that the incorporation of FixMatch is beneficial in the standard ST and ST++ framework. FixMatch training alongside the supervised training stage helps the model get a glimpse into the unlabeled data before actually training on them. In ST, the performance does not improve in as many cases as it does with ST++, primarily because it does not have separate stages to classify unlabeled images into reliable and unreliable ones. These stages in ST++ benefit tremendously from the FixMatch inclusion in the first stage, thus helping the model gain improvement in performance.
When incorporating FixMatch into the supervised training stage, there was a noticeable decrease in performance on the labeled images, as observed in Table 2. This is a commonly seen trade-off when we try to do two tasks in 1 stage. The performance dropped on supervised images is necessary for the model to perform well on unlabeled ones.
However, the performance in the later stage remains the same or even improves, suggesting that the trade-off in the first stage is necessary to improve the performance when extending the labeled images with unlabeled ones.
Figure 2: Sample of Ground truth mask, Non-FixMatch, and FixMatch mask generated from the dataset Zheng 1.
@@ 5.3 The quality of the pseudo-masks produced when incorporating FixMatch
The quality of the pseudo-masks between the two settings of FixMatch and non-FixMatch approaches varies. On the one hand, the pseudo-masks of easy, reliable images remain essentially the same between the two settings and on par with that of the ground truth mask, as illustrated in Figure 2. However, for more complicated, unreliable images, the masks generated can have a large deviation from the ground truth mask, as can be seen in Figure 3.
This suggests that incorporating FixMatch into the model has yet to be able to take on the nuances presented in challenging images. Therefore, there should be a more domain-specific way to account for these nuances.
@@
5.4 Generalization capability of the proposed approach
The proposed approach can be used in any dataset since the framework only includes general approaches commonly used in semi-supervised settings, namely self-training and FixMatch. However, the performance metrics are higher for datasets that are less complex compared to those that are more challenging. The performance of the proposed approach on Zheng 1 [9] and Zheng 2 [9] is impressive, while the performance on LISC [10] is low. This could be explained by LISC’s huge variability between its images 4. Therefore, the proposed approach has yet to resolve the intra-variation in challenging datasets and achieve consistent performance across different datasets.
Figure 4: LISC’s variability between its images leads to the model’s significantly reduced performances
6 Conclusion
In this paper, we propose a novel method for incorporating consistency regularization into standard and enhanced self-training approaches. In most cases, we find that this incorporation is beneficial. We also discover that there is a trade-off between excelling solely on labeled images and excelling on labeled images while using consistency regularization on unlabeled images. Furthermore, even with FixMatch included, the ST and ST++ approaches have yet to address the nuances of difficult images. To address these nuances, we plan to develop more domain-specific approaches in white blood cell segmentation in the future.
{credits}